find out the number of wave made by a bhor electron in one complete revolution in its 3rd orbit

Grade:11

1 Answers

Bhavya
6 years ago
Dear student
In general, the number of waves made by a Bohr electron in an orbit is equal to its quantum number.
According to Bohr’s postulate of angular momentum, in the 3rd orbit
Mur = n h/2π
Mur = 3 (h/2π) …..(i) [n = 3]
According to de Broglie relationship
λ = h/mu ….(ii)
Substituting (ii) in (i), we get
(h/λ) r = 3 (h/2π) or 3λ = 2πr
[∵ mu = h/λ]
Thus the circumference of the 3rd orbit is equal to 3 times the wavelength of electron i.e. the electron makesthreerevolution around the 3rd orbit.